OpenStudy (anonymous):

Factor completely. Remember to look first for a common factor and to check by multiplying. If a polynomial is prime, state this. -81p^2 + 18pq - q^2

OpenStudy (anonymous):

-81p^2 + 18pq - q^2 = -(9p^2) + 2(9p)(q) - q^2 You can prove that the polynomial is a perfect square, but for that, it needs to be like this: 81p^2 - 18pq + q^2. Are you sure that the question is right. If its right, then I think its prime.

OpenStudy (anonymous):

yes I am sure, that is the question, I have 2 more questions, This question is worked all the way through but I have no idea what was done to get the answer 1. Factor completely. If a polynomial is prime, state this 9t^2 + 14t + 5 t^2 + 7/9t = -5/9 +(7/9)^2 t^2 + 7/9t = (- 45 + 49)/81 (t + 7/9)^2 = 4/81 t + 7/9 = 2/9 t + 7/9 + 2/9 = t + 9/9 = t + 1 t + 7/9 - 2/9 = t + 5/9 = 9t + 5 final answer is (t+1)(9t+5) 2. If the area of a rectangle is given by 100g^2 - 220gw + 121w^2 square feet, what are possible expressions for the length and width?